# Multiplicity of powers of path ideals of a line graph

### v1

@@ 3. Multiplicity
(4) $\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}:x_{n-t}u_{n-t+1}},z)=\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s-1}}),z).$
If $n>2t$, by Lemma 3.7.(2), we have $(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})=(I_{t}(L_{n-t-1})^{s},x_{n-t})$. Since $n-t-1=(a-1)t+b-1$ with $b-1\geq 0$, it follows that
$\dim\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}}=n-t-1-(a-1)=n-a-t,$
and thus
(5) $\begin{split}\operatorname{HS}(\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})},z)=\frac{1}{(1-z)^{t}}\operatorname{HS}(\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}},z)\\ =\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}},z)\end{split}$
If $n<2t$, then $a=1$ and so
(6) $\operatorname{HS}(\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})},z)=\operatorname{HS}(\frac{S}{(x_{n-t})},z)=\frac{1}{(1-z)^{n-1}}=\frac{1}{(1-z)^{n-a}}$
If $n>2t$, substitute equations (2), (3), (4), and (5) into equation (1), we obtain
$\begin{split}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s}},z)=&z^{t}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s-1}},z)+\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s}},z)\\ &-z^{1+t}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s-1}},z)-z^{t}\mathbf{Q}(\frac{S_{n-t-1}}{I_{t}(L_{n-t-1})^{s}},z).\end{split}$
By substituting $z=1$ into the above equation and using the induction hypothesis, it follows that
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s})-\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s-1})$
$\qquad\qquad-\operatorname{mult}(S_{n-t-1}/I_{t}(L_{n-t-1})^{s})$
$=\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}+\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}$
$~{}~{}~{}~{}-\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}-\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}-\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}-\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b\\ a\end{pmatrix}-\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
If $n<2t$, by substituting equations (2), (3), (4), and (6) into equation (1), we obtain
$\begin{split}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s}},z)=&z^{t}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s-1}},z)+\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s}},z)\\ &-z^{1+t}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s-1}},z)-z^{t}\end{split}$
By substituting $z=1$ into the above equation and using the induction hypothesis, it follows that:
@@
$=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s})-\operatorname{mult}(S_{n-1}/I_{t}(L_{n-1})^{s-1})-1$
$=\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-b\\ 1\end{pmatrix}+\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-1$
$=\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-b\\ 0\end{pmatrix}-\begin{pmatrix}s-1\\ s-1\end{pmatrix}\begin{pmatrix}t-b\\ 0\end{pmatrix}$
$=\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b+1\\ 1\end{pmatrix}-\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b\\ 0\end{pmatrix}$
$=\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-b\\ 1\end{pmatrix}.$
Case 2: $b=0$. We may assume $a>1$, since the case that $a=1$, i.e, $n=t$ has been addressed. Put $A_{i}:=(I_{t}(L_{n-i})^{s},x_{n-t+1}\cdots x_{n-i+1})$ and $B_{i}:=(A_{i},x_{n-i+1})$ for $i=1,\ldots,t$. According to Lemma 3.8, we obtain the following $t-1$ short exact sequences:
@@
Since $A_{1}=(I_{t}(L_{n-1})^{s},u_{n-t+1})$, we can use the first short exact sequence used in the proof of the first case and the short exact sequences above to obtain
$\operatorname{HS}(S/I_{t}(L_{n})^{s},z)=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+\operatorname{HS}(S/A_{1},z)$
$=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+z\operatorname{HS}(S/A_{2},z)+\operatorname{HS}(S/B_{1},z)$
$=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+z^{2}\operatorname{HS}(S/A_{3},z)+\operatorname{HS}(S/B_{1},z)+z\operatorname{HS}(S/B_{2},z)$
=$\cdots$= ⋯
$=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+z^{t-1}\operatorname{HS}(S/A_{t},z)+\mathop{\sum}\limits_{i=1}^{t-1}z^{i-1}\operatorname{HS}(S/B_{i},z).$
Since $A_{t}=B_{t}$ by their definitions, it follows that
(7) $\operatorname{HS}(S/I_{t}(L_{n})^{s},z)=z^{t}\operatorname{HS}(S/I_{t}(L_{n})^{s-1},z)+\mathop{\sum}\limits_{i=1}^{t}z^{i-1}\operatorname{HS}(S/B_{i},z)$
Take any $i\in\{1,\ldots,t\}$. By Lemma 3.8, we see that $B_{i}=(I_{t}(L_{n-i})^{s},x_{n-i+1})$. Since $n-i=(a-1)t+t-i$, it follows $\dim S_{n-i}/I_{t}(L_{n-i})^{s}=n-i-a+1$, and so
(8) $\operatorname{HS}(S/B_{i},z)=\frac{1}{(1-z)^{i-1}}\operatorname{HS}(S_{n-i}/I_{t}(L_{n-i})^{s},z)=\frac{\mathbf{Q}(S_{n-i}/I_{t}(L_{n-i})^{s},z)}{(1-z)^{n-a}}$
Substitute equation (8) into equation (7), we have
(9) $\mathbf{Q}((S/I_{t}(L_{n})^{s},z)=z^{t}\mathbf{Q}((S/I_{t}(L_{n})^{s-1},z)+\mathop{\sum}\limits_{i=1}^{t}z^{i-1}\mathbf{Q}(S_{n-i}/I_{t}(L_{n-i})^{s},z)$
@@
$\operatorname{mult}(S/I_{t}(L_{n})^{s})$
$=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\mathop{\sum}\limits_{i=1}^{t}\operatorname{mult}(S_{n-i}/I_{t}(L_{n-i})^{s})$
$=\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}+\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\mathop{\sum}\limits_{i=1}^{t}\begin{pmatrix}a+i-2\\ a-1\end{pmatrix}$
$=\begin{pmatrix}s+a-2\\ s-2\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}+\begin{pmatrix}s+a-2\\ s-1\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}$
$=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-1\\ a\end{pmatrix}.$
@@
According to Theorem 3.9, it follows that the multiplicity of $S/I_{t}(L_{n})^{s}$ is a polynomial of degree $a$ in $s$.
Acknowledgment: This research is supported by NSFC (No. 11971338)

### v2

@@ 3. Multiplicity
$\dim\frac{S}{I_{t}(L_{n-1})^{s}:x_{n-t}u_{n-t+1}}=\dim\frac{S}{I_{t}(L_{n-1})^{s-1}}=n-\xi.$
If $n>2t$, by Lemma 3.7.(2), we have $(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})=(I_{t}(L_{n-t-1})^{s},x_{n-t})$. Since $n-t-1=(\xi-1)t+\theta-1$ with $\theta-1\geq 0$, it follows that
$\dim\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})}=\dim\frac{S}{(I_{t}(L_{n-t-1})^{s},x_{n-t})}=n-1-(\xi-1)=n-\xi.$
By Lemma 3.9 and using the induction hypothesis, it follows that
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S/I_{t}(L_{n-1})^{s})$
$\qquad\qquad\qquad\qquad-\operatorname{mult}(S/I_{t}(L_{n-1})^{s-1})-\operatorname{mult}(S/I_{t}(L_{n-t-1})^{s})$
$=$ $\begin{pmatrix}s+\xi-2\\ s-2\end{pmatrix}\begin{pmatrix}\xi+t-\theta-1\\ \xi\end{pmatrix}+\begin{pmatrix}s+\xi-1\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta\\ \xi\end{pmatrix}$
$~{}~{}~{}~{}-\begin{pmatrix}s+\xi-2\\ s-2\end{pmatrix}\begin{pmatrix}\xi+t-\theta\\ \xi\end{pmatrix}-\begin{pmatrix}s+\xi-2\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta-1\\ \xi-1\end{pmatrix}$
$=$ $\begin{pmatrix}s+\xi-1\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta\\ \xi\end{pmatrix}-\begin{pmatrix}s+\xi-2\\ s-2\end{pmatrix}\begin{pmatrix}\xi+t-\theta-1\\ \xi-1\end{pmatrix}$
$-\begin{pmatrix}s+\xi-2\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta-1\\ \xi-1\end{pmatrix}$
$=$ $\begin{pmatrix}s+\xi-1\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta\\ \xi\end{pmatrix}-\begin{pmatrix}s+\xi-1\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta-1\\ \xi-1\end{pmatrix}$
$=$ $\begin{pmatrix}s+\xi-1\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta-1\\ \xi\end{pmatrix}.$
If $t<n\leq 2t$, then $\xi=1$ and so
$\dim\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})}=\dim\frac{S}{(x_{n-t})}=n-\xi.$
Since $\operatorname{mult}(S/(x_{n-t}))=1$, by applying the induction hypothesis together with Lemma 3.9, we can deduce that:
@@
$=$ $\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S/I_{t}(L_{n-1})^{s})-\operatorname{mult}(S/I_{t}(L_{n-1})^{s-1})-1$
$=$ $\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-\theta\\ 1\end{pmatrix}+\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-\theta+1\\ 1\end{pmatrix}-\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-\theta+1\\ 1\end{pmatrix}-1$
$=$ $\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-\theta+1\\ 1\end{pmatrix}-\begin{pmatrix}s-1\\ s-2\end{pmatrix}\begin{pmatrix}t-\theta\\ 0\end{pmatrix}-\begin{pmatrix}s-1\\ s-1\end{pmatrix}\begin{pmatrix}t-\theta\\ 0\end{pmatrix}$
$=$ $\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-\theta+1\\ 1\end{pmatrix}-\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-\theta\\ 0\end{pmatrix}$
$=$ $\begin{pmatrix}s\\ s-1\end{pmatrix}\begin{pmatrix}t-\theta\\ 1\end{pmatrix}.$
Case 2: $\theta=0$, i.e., $n=\xi t$. We may assume $\xi>1$, since the case that $\xi=1$ has been addressed. As in Lemma 3.8, we put for $i=1,\ldots,t$,
$A_{i}:=(I_{t}(L_{n-i})^{s},x_{n-t+1}\cdots x_{n-i+1})\mbox{\quad and \quad}B_{i}:=(A_{i},x_{n-i+1}).$
We must compute the dimensions of these graded $S$-modules. Fix $i\in\{1,\ldots,t\}$. By Lemma 3.8, we see that $B_{i}=(I_{t}(L_{n-i})^{s},x_{n-i+1})$. Since we $n-i=(\xi-1)t+t-i$ with $0\leq t-i<t$, it follows from Lemma 3.2 that $\operatorname{height}(I_{t}(L_{n-i})^{s})=\xi-1$ and so $\operatorname{height}(B_{i})=\operatorname{height}(I_{t}(L_{n-i})^{s})+1=\xi$. We next show that $\operatorname{height}(A_{i})=\xi$. Since $\operatorname{height}(I_{t}(L_{n-i})^{s})=\xi-1$, it follows that $\operatorname{height}(A_{i})$ is either $\xi-1$ or $\xi$. Suppose on the contrary that $\operatorname{height}(A_{i})=\xi-1$. Then there must exist a minimal prime ideal, say $\operatorname{\mathfrak{p}}$, of $I_{t}(L_{n-i})^{s}$ with $\operatorname{height}(\operatorname{\mathfrak{p}})=\xi-1$ that contains $A_{i}$. However, this can not be the case. To see why, we assume without loss the generality that $\operatorname{\mathfrak{p}}=(x_{i_{1}},\ldots,x_{i_{\xi-1}})$ with $1\leq i_{1}<\cdots<i_{\xi-1}\leq n-i$. By Lemma 3.3, we have $(\xi-2)t+t-i+1\leq i_{\xi-1}\leq(\xi-1)t$. This implies $A_{i}\nsubseteq\operatorname{\mathfrak{p}}$, leading to contradiction. Therefore, we conclude that $\operatorname{height}(A_{i})=\xi$. As a result, we have
$\dim\frac{S}{A_{i}}=\dim\frac{S}{B_{i}}=n-\xi\mbox{ for all }i=1,\ldots,t.$
Since $A_{1}=(I_{t}(L_{n-1})^{s},u_{n-t+1})$, according to Lemma 3.9 and considering the short exact sequence (1), it follows that
(2) $\operatorname{mult}(S/I_{t}(L_{n})^{s})=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S/A_{1}).$
On the other hand, Lemma 3.8 asserts the existence of the following $t-1$ short exact sequences:
@@
In view of these short exact sequences together with (2), we obtain
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S/A_{1})$
$=$ $\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S/A_{2})+\operatorname{mult}(S/B_{1})$
$=$ $\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S/A_{3})+\operatorname{mult}(S/B_{1})+\operatorname{mult}(S/B_{2})$
$=$ $\cdots$
$=$ $\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\operatorname{mult}(S/A_{t})+\mathop{\sum}\limits_{i=1}^{t-1}\operatorname{mult}(S/B_{i}).$
Since $A_{t}=B_{t}$, it follows that
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\mathop{\sum}\limits_{i=1}^{t}\operatorname{mult}(S/B_{i})$
@@
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\operatorname{mult}(S/I_{t}(L_{n})^{s-1})+\mathop{\sum}\limits_{i=1}^{t}\operatorname{mult}(S/I_{t}(L_{n-i})^{s})$
$=$ $\begin{pmatrix}s+\xi-2\\ s-2\end{pmatrix}\begin{pmatrix}\xi+t-1\\ \xi\end{pmatrix}+\begin{pmatrix}s+\xi-2\\ s-1\end{pmatrix}\mathop{\sum}\limits_{i=1}^{t}\begin{pmatrix}\xi+i-2\\ \xi-1\end{pmatrix}$
$=$ $\begin{pmatrix}s+\xi-2\\ s-2\end{pmatrix}\begin{pmatrix}\xi+t-1\\ \xi\end{pmatrix}+\begin{pmatrix}s+\xi-2\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-1\\ \xi\end{pmatrix}$
$=$ $\begin{pmatrix}s+\xi-1\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-1\\ \xi\end{pmatrix}.$
This completes the proof.
@@
According to [14, Theorem 1.1], $\operatorname{mult}(S/I_{t}(L_{n})^{s})$ is of polynomial type of degree $\leq\xi$ in $s$. However, Theorem 3.10 reveals that $\operatorname{mult}(S/I_{t}(L_{n})^{s})$ is a polynomial of degree $\xi$ in $s$. Consequently, Theorem 3.10 serves a valuable complement to [14, Theorem 1.1].
Acknowledgment: This research is supported by NSFC (No. 11971338). We would express our sincere gratitude to the referee for his/her careful reading and numberous comments that improve the presentation of our paper greatly.